OpenStudy (jenniferjuice):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (osanseviero):

If it is parallel, it has the same slope

OpenStudy (azureilai):

m=(-8-7)/(-2-3) y=mx+b plug in (-2,3) in equation after you get slope to solve for b.

OpenStudy (osanseviero):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (jenniferjuice):

wait so i just find the slope of (-8-7) and (-2-3) ?

OpenStudy (azureilai):

yeah you divide change in y over change in x to get slope

OpenStudy (osanseviero):

And parallel lines have same slope

OpenStudy (jenniferjuice):

4/6?

OpenStudy (osanseviero):

 Example of parallel lines, same slope

OpenStudy (jenniferjuice):

4/6?

OpenStudy (osanseviero):

\[\frac{ -8-7 }{ -2-3 }=\frac{ 15 }{ 5 }=3\]

OpenStudy (jenniferjuice):

wouldnt it be -15/-5?

OpenStudy (azureilai):

it simplifies down to 3.

OpenStudy (jenniferjuice):

so m=3 ?

OpenStudy (azureilai):

yes

OpenStudy (jenniferjuice):

oh okay ..thanks :)

OpenStudy (osanseviero):

A negative number multiplied by a negative numer is positive